# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 3.4 Rational solutions of PII for $\alpha\in\mathbb{N}$
There are well-known exact rational solutions of (43)–(45), which will be important for our study. For a start, when $\alpha=1$, the solution of (43)–(45) is simply $F=1/\xi$. This scenario comes from $A_{0}=-2$, which happens to apply for the single-soliton solution. Here, the inner problem (43)–(45) is insufficient to describe the early time behaviour (as a singularity $\xi_{0}=0$ does not correspond to a moving singularity in the original variables) and higher-order terms are required. This case is considered in appendix B.4.
More generally, there are rational solutions of PII for all $\alpha\in\mathbb{N}$, all of which satisfy (43)–(45). Each of these solutions is of the form
$F=-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{Q_{\alpha-1}(\xi)}{Q_{\alpha}(\xi)}\right),\quad\alpha\in\mathbb{N},$
@@
$Q_{n+1}Q_{n-1}=\xi Q_{n}^{2}-4(Q_{n}Q_{n}^{\prime\prime}-Q_{n}^{\prime 2}),\quad Q_{0}=1,\quad Q_{1}=\xi.$
These $Q_{n}$ are referred to as Yablonskii-Vorob’ev polynomials [83, 86]. Using the recursion relation, we find
$Q_{2}=\xi^{3}+4,\quad Q_{3}=\xi^{6}+20\xi^{3}-80,\quad Q_{4}=\xi(\xi^{9}+60\xi^{6}+11200),\quad\ldots.$ (47)
@@
$\alpha=1:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{1}{\xi}\right)=\frac{1}{\xi},$ (48)
$\alpha=2:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{\xi}{\xi^{3}+4}\right)=-\frac{1}{\xi}+\frac{3\xi^{2}}{\xi^{3}+4},$ (49)
$\alpha=3:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{\xi^{3}+4}{\xi^{6}+20\xi^{3}-80}\right)=-\frac{3\xi^{2}}{\xi^{3}+4}+\frac{6\xi^{2}(\xi^{3}+10)}{\xi^{6}+20\xi^{3}-80},$ (50)
@@
$=$ $\,\frac{1}{\xi}-\frac{6\xi^{2}(\xi^{3}+10)}{\xi^{6}+20\xi^{3}-80}+\frac{9\xi^{5}(\xi^{3}+40)}{\xi^{9}+60\xi^{6}+11200}.$ (51)
We can easily check that each of these satisfies the far-field condition (44) with (45).
In terms of the inner problem for KdV, (27)–(28), applying the formula (31), these rational solutions correspond to
$\alpha=1:\quad$ $f_{0}=-\frac{2}{\xi^{2}},$ (52)
$\alpha=2:\quad$ $f_{0}=-\frac{6\xi(\xi^{3}-8)}{(\xi^{3}+4)^{2}},$ (53)
$\alpha=3:\quad$ $f_{0}=-\frac{12\xi(\xi^{9}+600\xi^{3}+1600)}{(\xi^{6}+20\xi^{3}-80)^{2}},$ (54)
$\alpha=4:\quad$ $f_{0}=-\frac{\xi^{18}+48\xi^{15}+2520\xi^{12}-78400\xi^{9}-1881600\xi^{6}+12544000}{\xi^{2}(\xi^{9}+60\xi^{6}+11200)^{2}}$ (55)
One conclusion is that the double poles of $f_{0}$ for each $\alpha\in\mathbb{N}$ are located at the zeros of the polynomials $Q_{\alpha}(\xi)$. These occur in an almost triangular structure [21, 22, 63]. Note that, while solutions $F$ of PII have simple poles of residue either +1+1 or 1-1, the application of (31) shows that only those with residue +1+1 correspond to double poles of $f_{0}$ (for simple poles with residue 1-1, say at $\xi=\bar{\xi}$, the function $f_{0}$ takes the value $\bar{\xi}/2$). This is why, for $\alpha\in\mathbb{N}$, there are $\alpha^{2}$ simple poles of $F$, but only ${\textstyle\frac{1}{2}}\alpha(\alpha+1)$ double poles of $f_{0}$.
For later, we shall check the result here for $\alpha=2$, which will be used to compare with the well-known 2-soliton solution of (1). We see above the poles of $f_{0}$ for $\alpha=2$ (the zeros of $Q_{2}=\xi^{3}+4$) are at
$\xi_{0}=-2^{2/3},\quad 2^{2/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right),\quad 2^{2/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right),$ (56)
where $2^{2/3}\approx 1.587$. Thus, when $\alpha=2$ (that is, $A_{0}=-6$), our small-time analysis (see (29)) suggests there are three singularities that move like
$s(t)\sim x_{0}-2^{2/3}(3t)^{1/3},\quad s(t)\sim x_{0}+2^{2/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\quad s(t)\sim x_{0}+2^{2/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}$ (57)
@@
Similarly, we shall later check our results here for $\alpha=3$ against the 3-soliton solution of (1). In this case, the poles of $f_{0}$ for $\alpha=3$ (the zeros of $Q_{3}=\xi^{6}+20\xi^{3}-80$) are at
@@
$(10+6\sqrt{5})^{1/3},\quad(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right),\quad(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right),$ (58)
where $(-10+6\sqrt{5})^{1/3}\approx 1.506$ and $10+6\sqrt{5}\approx 2.861$. Thus, for $\alpha=3$ ($A_{0}=-12$), our analysis predicts there are six singularities that emerge from each $x_{0}$, and they propagate like (see (29))
$s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}(3t)^{1/3},\quad s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},$ (59)
$s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\quad s(t)\sim x_{0}-(10+6\sqrt{5})^{1/3}(3t)^{1/3},$ (60)
$s(t)\sim x_{0}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\quad s(t)\sim x_{0}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},$ (61)
@@ 3.5 Numerical solutions of (43)–(45)
By adapting the code used to run numerical simulations of PII in Fornberg & Weideman [40] (see also [39, 41]), we are able to generate numerical solutions of (43)–(45) for a selection of values of the parameter $\alpha$. Some of these results are presented in figure 9. To obtain these results, we treat (43) as an initial-value problem with $F(L\mathrm{e}^{-\pi\mathrm{i}/3})$ determined by an optimal truncation of the far-field condition (44) using the recurrence relation in (41) to obtain as many terms as needed in (40), where $L$ is some moderately large value of $L$ ($L=10$, say). Using the algorithm of Fornberg & Weideman [40], we integrate from $\xi=L\mathrm{e}^{-\pi\mathrm{i}/3}$ towards $\xi=0$ in order to estimate $F(0)$ and $F^{\prime}(0)$. Once these estimates are obtained, Fornberg & Weideman’s algorithm is again used, now integrating along the entire square computational domain. In cases where $F(\xi)$ is singular at or near $\xi=0$, instead of integrating until $\xi=0$ we integrate until $|\xi|<h$ for some small $h$, and then use that stopping location as an initial value for evaluating the solution on the complete domain. Lastly, we note that if there a singularity on or near the ray $\xi=\rho\mathrm{e}^{-\pi\mathrm{i}/3}$, Fornberg & Weideman’s algorithm is able to adjust the path of integration automatically, thus no adjustment is needed to handle this situation.
@@
Figure 9: Phase portraits of solutions of PII, (43)–(45), computed for various values of $\alpha$, namely $0.25$, $0.5$, $0.75$ and $1$ in the first row, $1.25$, $1.5$, $1.75$, and $2$ in the second row, and $2.25$, $2.5$, $2.75$ and $3$ in the third row. The colour denotes the phase of $F$, with red denoting real and positive, yellow imaginary and positive, light blue real and negative, and dark blue negative and imaginary. Note the images for $\alpha=1$, $2$ and $3$ in this figure are generated using the exact solutions (48)–(50), while the rest are computed using the algorithms in Fornberg & Weideman [40]. In the latter case, we can see some numerical error in some images around $\mathrm{Re}(\xi)\approx 8$.
@@
Returning to figure 9, the most obvious distinction between some of these examples and others is that the solutions for $\alpha\in\mathbb{N}$ do not have a lattice of poles in the far field, as these are rational solutions with a finite number of poles. Note that the exact solutions for $\alpha=1$, $2$ and $3$ are given by (48)–(50), and the images in figure 9 are generated from these exact results (numerical solutions for $\alpha=1$, $2$ and $3$ are visually indistinguishable from the exact results for $7\lesssim\mathrm{Re}(\xi)\lesssim 7$, $7\lesssim\mathrm{Im}(\xi)\lesssim 7$, but show some numerical error outside of this region). Recall that our solutions of PII have simple poles with reside either +1+1 or 1-1. Using the solution for $\alpha=1$ in figure 9 (right panel of first row), namely $F=1/\xi$, as a reference, it should be clear how the colour scheme in the phase portrait displays a simple pole with residue +1+1 (with colouring red, dark blue, light blue, yellow, then red again, as we traverse once around the pole in the positive direction, starting at $\theta=0$). For example, we see the solution for $\alpha=2$ in figure 9 (right panel of second row), which has simple poles with residue +1+1 at the three points given by (56). The other pole for $\alpha=2$, located at the origin, has residue 1-1 (with colouring light blue, yellow, red, dark blue and light blue as we traverse around the pole starting at $\theta=0$). For $\alpha=3$ in figure 9 (right panel of third row), the six poles with residue +1+1 (58) form an almost triangular shape, while the three remaining poles with residue 1-1 lie inside this almost-triangle, and so on. Clear images of the pole fields for these and other rational solutions, together with the location of the simple zeros, can be found in Fornberg & Weideman [40], for example.
@@
For the solutions in figure 9 for $\alpha\notin\mathbb{N}$, the lattice of poles in the far field is restricted to a sector of angle $2\pi/3$, which is why these are called tritronquée solutions. The pole-free sector in the far field is $\pi<\mathrm{arg}(\xi)<\pi/3$. Another perspective of this lattice of poles is shown in figure 10, where there are a few analytic landscape plots, from which it is easy to appreciate where the poles are located. To help distinguish between poles with residue +1+1 or 1-1, we have added white and black dots to a phase portrait in figure 11. For this solution, drawn for $\alpha=5/2$, there are four poles (three with residue +1+1 and one with residue 1-1) that sit inside $\pi<\mathrm{arg}(\xi)<\pi/3$. As $\alpha$ increases, these “move” in the general direction $\pi/3$ and ultimately are located at known positions when $\alpha=3$ (the three with residue +1+1 are the first, third and sixth poles in equation (58)). Thus, we see that the property of a pole-free sector only holds in the far field, not the near field. To be clear, while in this section we are discussing solutions of (43)–(45), when we apply the change of variables (31), it is only the poles with residue +1+1 that are important for our solution $f_{0}$, and therefore it is only these poles that are relevant for the original KdV problem.
@@
Figure 10: Analytic landscape plots for numerical solutions of PII, (43)–(45), computed with $\alpha=0.5$, $1.5$ and $2.5$. The surfaces are $|F(\xi)|$, while the colour scheme is the same as in figure 9. The clear spikes in $|F(\xi)|$ correspond to simple poles.

### v2

@@ 3.5 Rational solutions of PII for $\alpha\in\mathbb{N}$
There are well-known exact rational solutions of (3.23)–(3.25), which will be important for our study. For a start, when $\alpha=1$, the solution of (3.23)–(3.25) is simply $F=1/\xi$. This scenario comes from $A_{0}=-2$, which happens to apply for the single-soliton solution. Here, the inner problem (3.23)–(3.25) is insufficient to describe the early time behaviour (as a singularity $\xi_{0}=0$ does not correspond to a moving singularity in the original variables) and higher-order terms are required. This case is considered in appendix B.4.
More generally, there are rational solutions of PII for all $\alpha\in\mathbb{N}$, all of which satisfy (3.23)–(3.25). Each of these solutions is of the form
$F=-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{Q_{\alpha-1}(\xi)}{Q_{\alpha}(\xi)}\right),\quad\alpha\in\mathbb{N},$ (3.27)
@@
$Q_{n+1}Q_{n-1}=\xi Q_{n}^{2}-4\left(Q_{n}\frac{\mathrm{d}^{2}Q_{n}}{\mathrm{d}\xi^{2}}-\left(\frac{\mathrm{d}Q_{n}}{\mathrm{d}\xi}\right)^{2}\right),\quad Q_{0}=1,\quad Q_{1}=\xi.$
These $Q_{n}$ are referred to as Yablonskii-Vorob’ev polynomials [94, 97]. They are special cases of the Adler-Moser polynomials, which are associated with rational solutions of the KdV equation [4, 5]; various other properties of the Yablonskii-Vorob’ev polynomials are summarised in Refs [10, 25, 63, 80, 87].
Using the recursion relation (or, equivalently, a determinant representation derived in [62]), we find
$Q_{2}=\xi^{3}+4,\quad Q_{3}=\xi^{6}+20\xi^{3}-80,\quad Q_{4}=\xi(\xi^{9}+60\xi^{6}+11200),\quad\ldots.$ (3.28)
@@
$\alpha=1:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{1}{\xi}\right)=\frac{1}{\xi},$ (3.29)
$\alpha=2:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{\xi}{\xi^{3}+4}\right)=-\frac{1}{\xi}+\frac{3\xi^{2}}{\xi^{3}+4},$ (3.30)
$\alpha=3:\quad F=$ $-\frac{\mathrm{d}}{\mathrm{d}\xi}\ln\left(\frac{\xi^{3}+4}{\xi^{6}+20\xi^{3}-80}\right)=-\frac{3\xi^{2}}{\xi^{3}+4}+\frac{6\xi^{2}(\xi^{3}+10)}{\xi^{6}+20\xi^{3}-80},$ (3.31)
@@
$=$ $\,\frac{1}{\xi}-\frac{6\xi^{2}(\xi^{3}+10)}{\xi^{6}+20\xi^{3}-80}+\frac{9\xi^{5}(\xi^{3}+40)}{\xi^{9}+60\xi^{6}+11200}.$ (3.32)
We can easily check that each of these satisfies the far-field condition (3.24) with (3.25). Further details of rational solutions of PII are provided in Refs [17, 73]
In terms of the inner problem for KdV, (3.5)–(3.6), applying the formula (3.9), these rational solutions correspond to
$\alpha=1:\quad$ $f_{0}=-\frac{2}{\xi^{2}},$ (3.33)
$\alpha=2:\quad$ $f_{0}=-\frac{6\xi(\xi^{3}-8)}{(\xi^{3}+4)^{2}},$ (3.34)
$\alpha=3:\quad$ $f_{0}=-\frac{12\xi(\xi^{9}+600\xi^{3}+1600)}{(\xi^{6}+20\xi^{3}-80)^{2}},$ (3.35)
$\alpha=4:\quad$ $f_{0}=-\frac{20(\xi^{18}+48\xi^{15}+2520\xi^{12}-78400\xi^{9}-1881600\xi^{6}+12544000)}{\xi^{2}(\xi^{9}+60\xi^{6}+11200)^{2}}.$ (3.36)
One conclusion is that the double poles of $f_{0}$ for each $\alpha\in\mathbb{N}$ are located at the zeros of the polynomials $Q_{\alpha}(\xi)$. These occur in an almost triangular structure [10, 17, 26, 27, 73]. Note that, while solutions $F$ of PII have simple poles of residue either +1+1 or 1-1, the application of (3.9) shows that only those with residue +1+1 correspond to double poles of $f_{0}$ (for simple poles with residue 1-1, say at $\xi=\bar{\xi}$, the function $f_{0}$ takes the value $\bar{\xi}/2$). This is why, for $\alpha\in\mathbb{N}$, there are $\alpha^{2}$ simple poles of $F$, but only ${\textstyle\frac{1}{2}}\alpha(\alpha+1)$ double poles of $f_{0}$.
A more direct way of deriving the rational solutions of $f_{0}$ is via
$f_{0}=2\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\ln Q_{\alpha}(\xi),\quad\alpha\in\mathbb{N}$
(see Clarkson [25], for example, where $f_{0}$ is denoted by $W$ in equation (46) in Clarkson’s paper). This formula has the advantage of requiring only one Yablonskii-Vorob’ev polynomial $Q_{n}$, whereas (3.27) involves two consecutive polynomials.
For later, we shall check the result here for $\alpha=2$, which will be used to compare with the well-known 2-soliton solution of (1.1). We see above that the poles of $f_{0}$ for $\alpha=2$ (the zeros of $Q_{2}=\xi^{3}+4$) are at
$\xi_{0}=-2^{2/3},\quad 2^{2/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right),\quad 2^{2/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right),$ (3.37)
where $2^{2/3}\approx 1.587$. Thus, when $\alpha=2$ (that is, $A_{0}=-6$), our small-time analysis (see (3.7)) suggests there are three singularities that move like
$s(t)\sim x_{0}-2^{2/3}(3t)^{1/3},\,\,s(t)\sim x_{0}+2^{2/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\,\,s(t)\sim x_{0}+2^{2/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}$ (3.38)
@@
Similarly, we shall later check our results here for $\alpha=3$ against the 3-soliton solution of (1.1). In this case, the poles of $f_{0}$ for $\alpha=3$ (the zeros of $Q_{3}=\xi^{6}+20\xi^{3}-80$) are at
@@
$(10+6\sqrt{5})^{1/3},\quad(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right),\quad(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right),$ (3.39)
where $(-10+6\sqrt{5})^{1/3}\approx 1.506$ and $10+6\sqrt{5}\approx 2.861$. Thus, for $\alpha=3$ ($A_{0}=-12$), our analysis predicts there are six singularities that emerge from each $x_{0}$, and they propagate like (see (3.7))
$s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}(3t)^{1/3},\quad s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},$ (3.40)
$s(t)\sim x_{0}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\quad s(t)\sim x_{0}-(10+6\sqrt{5})^{1/3}(3t)^{1/3},$ (3.41)
$s(t)\sim x_{0}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},\quad s(t)\sim x_{0}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3},$ (3.42)
@@
By adapting the code used to run numerical simulations of PII in Fornberg & Weideman [46] (referred to as the “pole field solver”; see also [45, 47]), we are able to generate numerical solutions of (3.23)–(3.25) for a selection of values of the parameter $\alpha$. Some of these results are presented in figure 9. To obtain these results, we treat (3.23) as an initial-value problem with $F(L\mathrm{e}^{-\pi\mathrm{i}/3})$ determined by an optimal truncation of the far-field condition (3.24) using the recurrence relation in (3.21) to obtain as many terms as needed in (3.20), where $L$ is some moderately large value of $L$ ($L=10$, say). Using the pole field solver [46], we integrate from $\xi=L\mathrm{e}^{-\pi\mathrm{i}/3}$ towards $\xi=0$ in order to estimate $F(0)$ and $F^{\prime}(0)$. Once these estimates are obtained, the pole field solver algorithm is again used, now integrating along the entire square computational domain. In cases where $F(\xi)$ is singular at or near $\xi=0$, instead of integrating until $\xi=0$ we integrate until $|\xi|<h$ for some small $h$, and then use that stopping location as an initial value for evaluating the solution on the complete domain. Lastly, we note that if there a singularity on or near the ray $\xi=\rho\mathrm{e}^{-\pi\mathrm{i}/3}$, the pole field solver algorithm is able to adjust the path of integration automatically, thus no adjustment is needed to handle this situation.
@@
Figure 9: Phase portraits of solutions of PII, (3.23)–(3.25), computed for various values of $\alpha$, namely $0.25$, $0.5$, $0.75$ and $1$ in the first row, $1.25$, $1.5$, $1.75$, and $2$ in the second row, and $2.25$, $2.5$, $2.75$ and $3$ in the third row. The colour denotes the phase of $F$, with red denoting real and positive, yellow imaginary and positive, light blue real and negative, and dark blue negative and imaginary. Note the images for $\alpha=1$, $2$ and $3$ in this figure are generated using the exact solutions (3.29)–(3.31), while the rest are computed using the pole field solver [46]. In the latter case, we can see some numerical error in some images around $\mathrm{Re}(\xi)\approx 8$.
@@
Returning to figure 9, the most obvious distinction between some of these examples and others is that the solutions for $\alpha\in\mathbb{N}$ do not have a lattice of poles in the far field, as these are rational solutions with a finite number of poles. Note that the exact solutions for $\alpha=1$, $2$ and $3$ are given by (3.29)–(3.31), and the images in figure 9 are generated from these exact results (numerical solutions for $\alpha=1$, $2$ and $3$ are visually indistinguishable from the exact results for $7\lesssim\mathrm{Re}(\xi)\lesssim 7$, $7\lesssim\mathrm{Im}(\xi)\lesssim 7$, but show some numerical error outside of this region). Recall that our solutions of PII have simple poles with reside either +1+1 or 1-1. Using the solution for $\alpha=1$ in figure 9 (right panel of first row), namely $F=1/\xi$, as a reference, it should be clear how the colour scheme in the phase portrait displays a simple pole with residue +1+1 (with colouring red, dark blue, light blue, yellow, then red again, as we traverse once around the pole in the positive direction, starting at $\theta=0$). For example, we see the solution for $\alpha=2$ in figure 9 (right panel of second row), which has simple poles with residue +1+1 at the three points given by (3.37). The other pole for $\alpha=2$, located at the origin, has residue 1-1 (with colouring light blue, yellow, red, dark blue and light blue as we traverse around the pole starting at $\theta=0$). For $\alpha=3$ in figure 9 (right panel of third row), the six poles with residue +1+1 (3.39) form an almost triangular shape, while the three remaining poles with residue 1-1 lie inside this almost-triangle, and so on. Clear images of the pole fields for these and other rational solutions, together with the location of the simple zeros, can be found in Fornberg & Weideman [46], for example.
@@
For the solutions in figure 9 for $\alpha\notin\mathbb{N}$, the lattice of poles in the far field is restricted to a sector of angle $2\pi/3$, which is why these are called tritronquée solutions. The pole-free sector in the far field is $\pi<\mathrm{arg}(\xi)<\pi/3$. Another perspective of this lattice of poles is shown in figure 10, where there are a few analytic landscape plots, from which it is easy to appreciate where the poles are located. To help distinguish between poles with residue +1+1 or 1-1, we have added white and black dots to a phase portrait in figure 11. For this solution, drawn for $\alpha=5/2$, there are four poles (three with residue +1+1 and one with residue 1-1) that sit inside $\pi<\mathrm{arg}(\xi)<\pi/3$. As $\alpha$ increases, these “move” in the general direction $\pi/3$ and ultimately are located at known positions when $\alpha=3$ (the three with residue +1+1 are the first, third and sixth poles in equation (3.39)). Thus, we see that the property of a pole-free sector only holds in the far field, not the near field. To be clear, while in this section we are discussing solutions of (3.23)–(3.25), when we apply the change of variables (3.9), it is only the poles with residue +1+1 that are important for our solution $f_{0}$, and therefore it is only these poles that are relevant for the original KdV problem.
@@
Figure 10: Analytic landscape plots for numerical solutions of PII, (3.23)–(3.25), computed with $\alpha=0.5$, $1.5$ and $2.5$. The surfaces are $|F(\xi)|$, while the colour scheme is the same as in figure 9. The clear spikes in $|F(\xi)|$ correspond to simple poles.